# AnyECG: Foundational Models for Electrocardiogram Analysis

### v1

@@ Abstract
Electrocardiogram (ECG), a non-invasive and affordable tool for cardiac monitoring, is highly sensitive in detecting acute heart attacks. However, due to the lengthy nature of ECG recordings, numerous machine learning methods have been developed for automated heart disease detection to reduce human workload. Despite these efforts, performance remains suboptimal. A key obstacle is the inherent complexity of ECG data, which includes heterogeneity (e.g., varying sampling rates), high levels of noise, demographic-related pattern shifts, and intricate rhythm-event associations. To overcome these challenges, this paper introduces AnyECG, a foundational model designed to extract robust representations from any real-world ECG data. Specifically, a tailored ECG Tokenizer encodes each fixed-duration ECG fragment into a token and, guided by proxy tasks, converts noisy, continuous ECG features into discrete, compact, and clinically meaningful local rhythm codes. These codes encapsulate basic morphological, frequency, and demographic information (e.g., sex), effectively mitigating signal noise. We further pre-train the AnyECG to learn rhythmic pattern associations across ECG tokens, enabling the capture of cardiac event semantics. By being jointly pre-trained on diverse ECG data sources, AnyECG is capable of generalizing across a wide range of downstream tasks where ECG signals are recorded from various devices and scenarios. Experimental results in anomaly detection, arrhythmia detection, corrupted lead generation, and ultra-long ECG signal analysis demonstrate that AnyECG learns common ECG knowledge from data and significantly outperforms cutting-edge methods in each respective task.
Keywords ECG representation, Cardiac Diagnosis, Denoising Reconstruction, Foundation Model
1 Introduction
@@
Sequence models, such as large language models, developed using large-scale data in the wild have shown significant advantages in learning robust representations and demonstrated robustness in downstream tasks. However, adapting sequence model pre-training approaches to ECG data in real-world settings poses unique challenges: (1) Heterogeneity: real-world ECG signals vary in length, sampling rate, and the number of channels due to differences in devices and scenarios. A unified model (e.g., with fixed tokenizer settings) is needed to effectively manage this diversity while maintaining temporal resolution and avoiding the introduction of artifacts. (2) Low SNR: ECG signals inherently have a low SNR, and pathological waveforms are often subtle, making it easy for noise to interfere with the understanding of critical features. (3) Demographic shift: ECG waveforms can vary due to patient demographics (e.g., age, sex, ethnicity). For instance, pediatric ECGs differ from adults in disease presentation and heart rate [5], and distinct ethnic groups may exhibit unique ECG characteristics [6], which hinder models from generalizing across diverse populations. (4) Implicit rhythm-event association: The systematic arrangement of rhythm patterns may indicate some cardiac events. ECG analysis depends on identifying these rhythm associations and event patterns. However, noisy real-world ECG data make it difficult for models to capture sequential relationships, hindering the accurate cardiac event detection.
@@
To overcome these challenges, we introduce AnyECGs, a family of ECG foundational models designed for robust representation learning on ECG signals in diverse, real-world settings. The development of AnyECG involves two main pre-training phases: the ECG Tokenizer pre-training and the entire AnyECG foundation model pre-training. The first phase captures key local rhythmic patterns from noisy ECG signals, while the second learns associations across the rhythmic patterns that implies cardiac events. Specifically, we first pad any ECG signals to unified length, segment a signal into a collection of fixed-duration fragments, project each fragment into a token orderly, and pad missing channels, standardizing the diversity in sampling rates, lengths, and channel numbers. We further design a new hierarchical modeling approach to tackle ultra-long ECG signal (solving challenge (1)). In the ECG Tokenizer pre-training phase, a Rhythm Codebook is established to capture the key local morphological and frequency features inherent in ECG signals. The ECG Tokenizer extracts ECG features that are closely aligned with these Rhythm Codes, effectively reducing noise by matching the input patches to those representative codes (solving challenge (2)). Additionally, the extracted ECG features are also required to recover demographic information about the patient (addressing challenge (3)). Then, we apply ‘masked modeling’ approach in the AnyECG pretraining phase, where the model predicts Rhythm Code indices to fill in masked patches. This approach encourages the recovery of masked ECG patches based on their relationship with unmasked ECG patches, enabling the model to learn cardiac event semantics that are essential for downstream tasks, thus addressing challenge (4). With these designs, AnyECG can facilitate knowledge transfer across various ECG sources in the wild, enabling to learn shared ECG and cardiac event knowledge that is applicable to downstream tasks. Our contributions are listed below.
  • ECG Foundational Model: We introduce AnyECG, a foundational ECG model that unifies representation learning by capturing important local rhythm patterns in ECG signals and their semantic relationships, providing a flexible framework adaptable to any ECG signals for various downstream applications.
  • ECG Tokenizer: We present an ECG tokenizer that extracts compact, noise-resilient Rhythm Codes utilizing the Multi-View Synergistic Decoder that reconstructs these codes from Morphology, Frequency, and Demography perspectives to capture essential diagnostic features and improve generalization across diverse populations.
@@
  • Various Downstream Tasks Adaptability: By pretrained to learn the associations among ECG tokens, our AnyECG is sensitive to potential cardiac events, demonstrating strong generalization capabilities across various downstream tasks, including anomaly detection, arrhythmia detection, corrupted lead generation, and ultra-long ECG signal analysis.
2 Methodology
@@
Figure 1: Overall architecture and pre-training pipeline of AnyECG. AnyECG is pre-trained in two steps. The Patient Attribute Tokenizer is pre-trained through proxy tasks to embed morphology, frequency, and demography into tokens (up). Then, the entire AnyECG, along with the Patient Attribute Tokenizer, is further pre-trained by predicting the code indices of the masked tokens to learn the semantic relationships between tokens (bottom-left). The Cross-Masking approach restricts interactions of patches from the same lead or from the same position across different leads (bottom-right). LN: LayerNorm, Conv: 1D convolution with kernel size of 15.
Our proposed AnyECG adopts a common Transformer architecture, incorporating a novel attention module and a special tokenizer that can be adapted to both the self-supervised learning pretraining pipeline and various downstream tasks for any ECG signals, as shown in Figure 1. The self-supervised learning process of AnyECG is divided into two phases: pretraining the ECG Tokenizer and pretraining the AnyECG backbone.
2.1 AnyECG Architecture
ECG Signal Pre-Processing.
@@ 2.1 AnyECG Architecture
ECG Tokenizer.
The objective of ECG Tokenizer is to effectively capture both the temporal and spatial features of ECG signals and generate an embedding $H\in\mathbb{R}^{N\times d}$ from $X^{\prime}$. We firstly use a temporal encoder to learn local temporal patterns by independently processing each ECG patch $x_{j,k}$. The temporal encoder consists of one 1D convolution, a group normalization, and a GELU activation function. Then, a spatial encoder with 4 Transformer blocks is used. Each patch $x_{j,k}$ is processed by the temporal encoder and spatial encoder sequentially to obtain an embedding $h^{\prime}_{j,k}\in\mathbb{R}^{d}$. To enhance the model’s understanding of the temporal sequence and leads relationships, learnable temporal position encoding $\tau_{k}\in\mathbb{R}^{d}$ and lead position encoding $\sigma_{j}\in\mathbb{R}^{d}$ are along the temporal dimension $k$ and channel dimension $j$, separately. The output of the ECG Tokenizer is the temporal-spatial encoding ${h}_{j,k}=h^{\prime}_{j,k}+\tau_{k}+\sigma_{j}$.
Cross-Mask Attention (CMA).
Unlike other sequential data like text, ECG signals typically include multiple leads, with the signals at the same positions across leads providing complementary information [8]. Therefore, in contrast to conventional multi-head self-attention, we introduce CMA, which differentiates the structure of our AnyECG. CMA allows each patch to interact only with patches within relevant channels (i.e. leads) and temporal contexts. We apply CMA as the attention module within the Transformer blocks of both the ECG Tokenizer and AnyECG backbone. Given input $H$, $Q=\text{LayerNorm}({H})W_{Q}$, $K=\text{LayerNorm}({H})W_{K}$, $V={H}W_{V}$, where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{\text{model}}}$ are learnable projection matrices, $\text{LayerNorm}(\cdot)$ denotes layer normalization, and $d_{\text{model}}$ is the model dimension. The CMA is computed as:
$\text{CMA}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}+M}{\sqrt{d_{\text{head}}}}\right)V\end{aligned},\ \ \ \ \begin{aligned} M_{i,j}=\begin{cases}0,&\text{if }j\in\mathcal{A}(i)\\ -\infty,&\text{otherwise}\end{cases}\end{aligned}$ (1)
@@
2.2 ECG Tokenizer Pretraining
2.2.1 ECG Tokenizer with Rhythm Codebook
@@
$i^{\ast}=\mathop{\mathrm{arg\,min}}_{i\in\{1,2,\dotsc,K\}}\left\|\frac{h_{j,k}}{\|h_{j,k}\|_{2}}-\frac{v_{i}}{\|v_{i}\|_{2}}\right\|^{2}$ (2)
@@
2.2.2 Multi-View Synergistic Decoder
@@
$\mathcal{L}_{\text{codebook}}=\sum_{j=1}^{C}\sum_{k=1}^{P}\left\|\operatorname{sg}\left(h_{j,k}\right)-v_{z_{j,k}}\right\|_{2}^{2}\end{aligned}\ \ \ \ \ \begin{aligned} \mathcal{L}_{\text{commitment}}=\beta\sum_{j=1}^{C}\sum_{k=1}^{P}\left\|h_{j,k}-\operatorname{sg}\left(v_{z_{j,k}}\right)\right\|_{2}^{2}\end{aligned}$ (7)
@@
$\mathcal{L}_{T}=\mathcal{L}_{\text{morphology}}+\mathcal{L}_{\text{frequency}}+\mathcal{L}_{\text{demography}}+\mathcal{L}_{\text{codebook}}+\mathcal{L}_{\text{commitment}}$ (8)
This loss function requires the reconstruction of both the temporal and frequency components of the ECG signal, while also ensuring the recovery of patient-specific factors for personalized modeling. Experiments in Appendix 7.4 shows the importance of each component in the total loss function.
2.3 AnyECG Masked Pre-training
@@
The masked pretraining facilitates the model in learning generic representations from the input data by capturing the implicit rhythm-event associations and sequential relationships crucial for ECG analysis, thereby enhancing its ability to capture the underlying cardiac event patterns in the ECG signals.
3 Downstream Application
@@
This section evaluates AnyECG’s performance across multiple ECG datasets to prove its generality. In Section 3.1, we summarize datasets utilized in the experiments. Section 3.2 explains the experimental setup in detail. In Section 3.3, we present the results of our experiments, benchmarking AnyECG against state-of-the-art methods across multiple tasks, including anomaly detection, arrhythmia detection, ECG lead generation, and ultra-long ECG sequence recognition. In Section 7.4 and 7.3, we also present ablation studies on hyperparameter selection and the necessity of two-stage pre-training.
3.1 ECG Datasets

### v2

# AnyECG: Foundational Models for Multitask Cardiac Analysis in Real-World Settings

@@ Abstract
Electrocardiogram (ECG), a non-invasive and affordable tool for cardiac monitoring, is highly sensitive in detecting acute heart attacks. However, due to the lengthy nature of ECG recordings, numerous machine learning methods have been developed for automated heart disease detection to reduce human workload. Despite these efforts, performance remains suboptimal. A key obstacle is the inherent complexity of ECG data, which includes heterogeneity (e.g., varying sampling rates), high levels of noise, demographic-related pattern shifts, and intricate rhythm-event associations. To overcome these challenges, this paper introduces AnyECG, a foundational model designed to extract robust representations from any real-world ECG data. Specifically, a tailored ECG Tokenizer encodes each fixed-duration ECG fragment into a token and, guided by proxy tasks, converts noisy, continuous ECG features into discrete, compact, and clinically meaningful local rhythm codes. These codes encapsulate basic morphological, frequency, and demographic information (e.g., sex), effectively mitigating signal noise. We further pre-train the AnyECG to learn rhythmic pattern associations across ECG tokens, enabling the capture of cardiac event semantics. By being jointly pre-trained on diverse ECG data sources, AnyECG is capable of generalizing across a wide range of downstream tasks where ECG signals are recorded from various devices and scenarios. The experimental results show that AnyECG achieves an average performance improvement of 6% across four critical tasks—anomaly detection, arrhythmia classification, corrupted lead generation, and ultra-long ECG recognition. AnyECG learns common ECG rhythm from data and significantly outperforms state-of-the-art methods in each of these tasks.
Index Terms:
ECG representation, Cardiac Diagnosis, Denoising Reconstruction, Foundation Model
I Introduction
@@
Sequence models, such as large language models, developed using large-scale data in the wild have shown significant advantages in learning robust representations and demonstrated robustness in downstream tasks. However, adapting sequence model pre-training approaches to ECG data in real-world settings poses unique challenges: (1) Heterogeneity: Real-world ECG signals exhibit significant variations in length (seconds to hours), sampling rates (100–1000 Hz), and channel numbers (1–12 leads) across acquisition devices. Conventional approaches struggle to reconcile these differences: sampling rate discrepancies fundamentally distort temporal patterns (e.g., a heartbeat spans 4$\times$ fewer samples at 125 Hz vs. 500 Hz), while unifying high-resolution long-term signals (e.g., >86M samples for 24-hour 1000 Hz data) exceeds computational limits of standard architectures. Existing methods either ignore these critical variations or impose aggressive downsampling that degrades diagnostically vital features. (2) Low SNR: ECG signals inherently have a low SNR, and pathological waveforms are often subtle, making it easy for noise to interfere with the understanding of critical features. Meanwhile, during signal extraction, deep learning models tend to prioritize the extraction of high-frequency signals (such as the QRS complex) while neglecting low-frequency features (such as the ST-T segment), leading to an imbalance in the model’s spectral sensitivity.
(3) Demographic shift: ECG waveforms can vary due to patient demographics (e.g., age, sex, ethnicity). For instance, pediatric ECGs differ from adults in disease presentation and heart rate [5], and distinct ethnic groups may exhibit unique ECG characteristics [6], which hinder models from generalizing across diverse populations.
@@
To overcome these challenges, we introduce AnyECGs, a family of ECG foundational models designed for robust representation learning on ECG signals in diverse, real-world settings. The development of AnyECG involves two main pre-training phases: the Rhythm Quantizer pre-training and the entire AnyECG foundation model pre-training. The first phase captures key local rhythmic patterns from noisy ECG signals, while the second learns associations across the rhythmic patterns that implies cardiac events. Specifically, we first pad any ECG signals to unified length, segment a signal into a collection of fixed-duration fragments, project each fragment into a token orderly, and pad missing channels, standardizing the diversity in sampling rates, lengths, and channel numbers. We further design a new hierarchical modeling approach to tackle ultra-long ECG signals. In the Rhythm Quantizer training phase, a Rhythm Codebook is established to capture the key local morphological and frequency features inherent in ECG signals. The ECG Tokenizer extracts ECG features that are closely aligned with these Rhythm Codes, effectively reducing noise by matching the input patches to those representative codes. Additionally, the extracted ECG features are also required to recover demographic information about the patient. Then, we apply ‘masked modeling’ approach in the AnyECG pre-training phase, where the model predicts Rhythm Code indices to fill in masked patches. This approach encourages the recovery of masked ECG patches based on their relationship with unmasked ECG patches, enabling the model to learn cardiac event semantics that are essential for downstream tasks. With these designs, AnyECG can facilitate knowledge transfer across various ECG sources in the wild, enabling to learn shared ECG and cardiac event knowledge that is applicable to downstream tasks. Our contributions are listed below.
  • Cardio-Sparse Attention:Inspired by physicians’ prioritization of key cardiac events, we design Cardio-Sparse Attention (CSA), a mechanism that focuses on critical rhythm phases while masking non-informative segments. By mimicking clinicians’ diagnostic focus on actionable patterns and ignoring redundant noise, this design reduces computational costs, enabling efficient analysis of ultra-long ECG recordings.
  • Rhythm Quantizer: We propose a Rhythm Quantizer framework that transforms raw ECG signals into discrete, noise-resilient Rhythm Codes through a multi-view decoding process. By jointly reconstructing morphological waveforms (e.g., QRS complexes), time-frequency patterns (via wavelet decomposition), and demographic attributes (e.g., age-related T-wave variations), this framework explicitly separates pathological features from artifacts. The generated codes provide a unified representation for heterogeneous ECG data, enabling robust analysis in low-SNR scenarios.
@@
  • Foundation Model for Unified Cardiac Diagnostics: By pretrained to learn the associations among ECG tokens, our AnyECG is sensitive to potential cardiac events and achieves state-of-the-art performance including anomaly detection, arrhythmia detection, corrupted lead generation, and ultra-long ECG signal analysis. The two-stage pre-training paradigm ensures adaptability to fragmented clinical workflows, providing a unified solution for real-world ECG diagnostics.
II Related Works
Heterogeneous ECG Signal Analysis and Classification
Self-supervised ECG Representation Learning
III Methodology
@@
Figure 1: Overall architecture and pre-training pipeline of AnyECG. AnyECG is pre-trained in two steps. The Rhythm Quantizer is pre-trained through proxy tasks to embed morphology, frequency, and demography into tokens (up). Then, the entire AnyECG, along with the ECG Tokenizer, is further pre-trained by predicting the code indices of the masked tokens to learn the semantic relationships between tokens (bottom-left). The Cardio-Sparse approach restricts interactions of patches from the same lead or from the same position across different leads (bottom-right). LN: LayerNorm, Conv: 1D convolution with kernel size of 15.
This section presents our proposed AnyECG, which is designed to be adaptable to both the self-supervised learning pre-training pipeline and various downstream tasks for any ECG signals, as illustrated in Figure 1. In Section III-A, we introduce the architecture of AnyECG, detailing its components and design principles. In Sections III-B and III-C, we describe the Rhythm Quantizer and the pre-training process of AnyECG. Finally, in Section III-D, we summarize the key takeaways.
III-A AnyECG Architecture
ECG Signal Pre-Processing
@@ III-A AnyECG Architecture
Rhythm Quantizer
The objective of Rhythm Quantizer is to effectively capture both the temporal and spatial features of ECG signals and generate an embedding $H\in\mathbb{R}^{N\times d}$ from $X^{\prime}$. We firstly use a temporal encoder to learn local temporal patterns by independently processing each ECG patch $x_{j,k}$. The temporal encoder consists of one 1D convolution, a group normalization, and a GELU activation function. Then, a spatial encoder with 4 Transformer blocks is used. Each patch $x_{j,k}$ is processed by the temporal encoder and spatial encoder sequentially to obtain an embedding $h^{\prime}_{j,k}\in\mathbb{R}^{d}$. To enhance the model’s understanding of the temporal sequence and leads relationships, learnable temporal position encoding $\tau_{k}\in\mathbb{R}^{d}$ and lead position encoding $\sigma_{j}\in\mathbb{R}^{d}$ are along the temporal dimension $k$ and channel dimension $j$, separately. The output of the ECG Tokenizer is the temporal-spatial encoding ${h}_{j,k}=h^{\prime}_{j,k}+\tau_{k}+\sigma_{j}$.
Cardio-Sparse Attention (CSA)
Unlike other sequential data like text, ECG signals typically include multiple leads, with the signals at the same positions across leads providing complementary information [26]. Therefore, in contrast to conventional multi-head self-attention, we introduce CSA, which differentiates the structure of our AnyECG. CSA allows each patch to interact only with patches within relevant channels (i.e. leads) and temporal contexts. We apply CSA as the attention module within the Transformer blocks of both the ECG Tokenizer and AnyECG backbone. Given input $H$, $Q=\text{LayerNorm}({H})W_{Q}$, $K=\text{LayerNorm}({H})W_{K}$, $V={H}W_{V}$, where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{\text{model}}}$ are learnable projection matrices, $\text{LayerNorm}(\cdot)$ denotes layer normalization, and $d_{\text{model}}$ is the model dimension. The CSA is computed as:
$\text{CSA}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}+M}{\sqrt{d_{\text{head}}}}\right)V,$ (1)
$M_{i,j}$ =$=\begin{cases}0,&\text{if }j\in\mathcal{A}(i),\\ -\infty,&\text{otherwise}.\end{cases}$
@@
III-B Rhythm Quantizer Pre-training
III-B1 ECG Tokenizer with Rhythm Codebook
@@
$i^{\ast}=\mathop{\arg\min}_{i\in\{1,2,\dotsc,K\}}\left\|\frac{h_{j,k}}{\|h_{j,k}\|_{2}}-\frac{v_{i}}{\|v_{i}\|_{2}}\right\|^{2}$ (2)
@@
III-B2 Multi-View Synergistic Decoder
@@
$\mathcal{L}_{\text{codebook}}$ $=\sum_{j=1}^{C}\sum_{k=1}^{P}\left\|\operatorname{sg}\left(h_{j,k}\right)-v_{z_{j,k}}\right\|_{2}^{2},$ (7)
$\mathcal{L}_{\text{commitment}}$ $=\beta\sum_{j=1}^{C}\sum_{k=1}^{P}\left\|h_{j,k}-\operatorname{sg}\left(v_{z_{j,k}}\right)\right\|_{2}^{2}.$
@@
$\begin{split}\mathcal{L}_{T}=&\ \mathcal{L}_{\text{morphology}}+\mathcal{L}_{\text{frequency}}+\mathcal{L}_{\text{demography}}+\\ &\mathcal{L}_{\text{codebook}}+\mathcal{L}_{\text{commitment}}\end{split}$ (8)
This loss function requires the reconstruction of both the temporal and frequency components of the ECG signal, while also ensuring the recovery of patient-specific factors for personalized modeling. Experiments in Appendix IV-F shows the importance of each component in the total loss function.
III-C AnyECG Masked Pre-training
@@
The masked pre-training facilitates the model in learning generic representations from the input data by capturing the implicit rhythm-event associations and sequential relationships crucial for ECG analysis, thereby enhancing its ability to capture the underlying cardiac event patterns in the ECG signals.
III-D Takeaway
Our proposed AnyECG framework innovatively addresses three core challenges in current ECG Foundation Models, with the following key insights:
Heterogeneity in ECG signals—arising from varying sam-pling rates, lead configurations, and recording lengths—poses significant challenges for consistent analysis. To address this, standardized preprocessing is employed to unify signal resolution, ensuring compatibility across diverse devices by resampling signals to 300 Hz and segmenting them into fixed-duration patches. However, this preprocessing step introduces a side effect: the generation of long ECG patches, which can lead to computational inefficiency. To resolve this issue, CSA is introduced, which selectively focuses on critical rhythm phases and cross-lead dependencies, mimicking the diagnostic focus of clinicians. This sparse attention mechanism reduces redundant computations on non-informative segments, enabling efficient analysis of ultra-long sequences without sacrificing temporal resolution—a key limitation in prior work.
Low SNR, which obscures subtle pathological features, is mitigated by the Rhythm Quantizer. This module distills ECG signals into discrete tokens through three reconstruction objectives: (1) morphological reconstruction of disease-specific waveforms, (2) time-frequency alignment via wavelet decomposition to isolate rhythm irregularities (3) demographic attribute recovery (age-related T-wave variations). By explicitly encoding cardiac rhythms, the quantizer disentangles clinically actionable patterns from interference.
Demographic shifts—such as age, sex, and ethnicity biases—are tackled via a two-stage pre-training paradigm. First, local rhythm codes are learned to capture demographic-agnostic patterns. Second, masked modeling of global cardiac semantics links rhythm events to clinical contexts, ensuring robustness across populations. The model decouples age-related variations from pathological signatures. This approach bridges domain gaps between hospital-grade and wearable ECG data.
CSA and Rhythm Quantizer jointly address heterogeneity and noise, while two-stage pre-training ensures demographic invariance. The unified design establishes a foundation for scalable, device-agnostic cardiac diagnostics.
IV Downstream Application
@@
This section evaluates AnyECG’s performance across multiple ECG datasets to prove its generality. As illustrated in Figure 2, AnyECG is fine-tuned on heterogeneous ECG data to address four critical tasks: anomaly detection, arrhythmia classification, corrupted lead generation, and ultra-long ECG recognition. In Section IV-A, we summarize datasets utilized in the experiments. Section IV-C explains the experimental setup in detail. In Section IV-D, we present the results of our experiments, benchmarking AnyECG against state-of-the-art methods across multiple tasks, including anomaly detection, arrhythmia detection, ECG lead generation, and ultra-long ECG sequence recognition. In Section IV-F and IV-E, we also present ablation studies on hyperparameter selection and the necessity of two-stage pre-training.
Figure 2: Downstream Task of the ECG Foundation Model. The framework demonstrates fine-tuning AnyECG on heterogeneous ECG datasets, including portable devices, Holter monitors, standard 12-lead recordings, and Event monitor, to address four critical tasks: (1) Anomaly Detection (binary classification of normal vs. abnormal rhythms), (2) Arrhythmia Classification, (3) Ultra-Long ECG Analysis (continuous monitoring for ultra-long ECG event detection), and (4) Corrupted Lead Generation. AnyECG achieves task-specific adaptability through fine-tuning while preserving pre-learned cardiac semantics.
IV-A ECG Datasets